Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A bag contains a total of four billiard balls. The billiard

Expert replies
by BTGmoderatorAT » Thu Dec 14, 2017 6:03 am
A bag contains a total of four billiard balls. The billiard balls are numbered as follows: 2, 3, 5 and 7.

Scenario 1: Two billiard balls are pulled out of the bag. What is the probability that the sum of the numbers appearing on the balls is odd?
Scenario 2: A billiard ball is pulled out of the bag then placed back in. Another billiard ball is pulled out of the bag (which may or may not be identical to the first) then placed back in. What is the probability that the sum of the two balls is odd?

I'm confused how to set up the formulas here. Can any experts help?
Join the discussion
Source: — Problem Solving |

by Jay@ManhattanReview » Fri Dec 15, 2017 12:41 am
ardz24 wrote:A bag contains a total of four billiard balls. The billiard balls are numbered as follows: 2, 3, 5 and 7.

Scenario 1: Two billiard balls are pulled out of the bag. What is the probability that the sum of the numbers appearing on the balls is odd?
Scenario 2: A billiard ball is pulled out of the bag then placed back in. Another billiard ball is pulled out of the bag (which may or may not be identical to the first) then placed back in. What is the probability that the sum of the two balls is odd?

I'm confused how to set up the formulas here. Can any experts help?
Scenario 1: Two billiard balls are pulled out of the bag. What is the probability that the sum of the numbers appearing on the balls is odd?

A sum of odd can be formed if a ball numbered 2 is drawn. So, the samples are (2, 3); (2 5); (2, 7): a total 3 samples out of 4C2 = 4.3/1.2 = 6 samples.

Probability = 3/6 = 1/2

Scenario 2: A billiard ball is pulled out of the bag then placed back in. Another billiard ball is pulled out of the bag (which may or may not be identical to the first) then placed back in. What is the probability that the sum of the two balls is odd?

Case 1: A ball number 2 is drawn and one ball from (3, 5, 7) is drawn

Probability = (Probability of drawing a numbered 2) * (Probability of drawing a numbered 3, 5 or 7) = (1/4) * (3/4) = 3/16

Case 2: A ball from (3, 5, 7) is drawn and then a ball number 2 is drawn

Probability = (Probability of drawing a numbered 3, 5 or 7) * (Probability of drawing a numbered 2) = (3/4) * (1/4) = 3/16

Total probability = 3/16 + 3/18 = 3/8.

Hope this helps!

-Jay
_________________
Manhattan Review GMAT Prep

Locations: Pasadena | Warsaw | Cairo | Kuala Lumpur | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion